# Symmetry Analysis of the Non-Hermitian Electro-Optic Effect in Crystals

### v1

@@ Abstract
We investigate how crystal symmetry tailors the non-Hermitian electro-optic effect driven by the Berry curvature dipole. Specifically, we demonstrate the critical influence of the material’s point group symmetry and external electric biases in shaping this effect, leading to current-induced optical gain and non-reciprocal optical responses. Through a symmetry-based analysis of the crystallographic point groups, we identify how different symmetries affect the electro-optic response, enabling the engineering of polarization-dependent optical gain without the need for gyrotropic effects. In particular, we demonstrate that the non-Hermitian electro-optic response in a broad class of crystals is characterized by linear dichroic gain. In this type of response, the eigenpolarizations that activate the gain or dissipation are linearly polarized. Depending on the specific symmetry point group, it is possible to achieve gain (or dissipation) for all eigenpolarizations or to observe polarization-dependent gain and dissipation. Weyl semimetals emerge as promising candidates for realizing significant non-Hermitian electro-optic effects and linear dichroic gain. We further examine practical applications by studying the reflectance of biased materials in setups involving mirrors, demonstrating how optical gain and attenuation can be controlled via symmetry and bias configurations.
@@ I Introduction
Several studies have further linked the BD to optical effects, including SHG, the kinetic Faraday effect, where the rotatory power is proportional to the electric current and reverses sign with the applied electric field Vorob’ev et al. (1979); Tsirkin et al. (2018); König et al. (2019), and the circular photogalvanic effect, where the photocurrent reverses sign with the helicity of light Ivchenko and Pikus (1978); Asnin et al. (1978); Deyo et al. (2009); Xu et al. (2018). These effects are particularly prominent in materials with small band gaps, where the Berry curvature gets more enhanced and significantly influences interband transitions. Crucially, many second-order optical responses have been shown to directly correlate with the Berry curvature dipole, establishing them as effective probes for investigating a material’s electronic structure and symmetry Xu et al. (2018); Bhalla et al. (2022).
Recent proposals have demonstrated that electro-optic effects can lead to non-reciprocal optical gain, where the amplification of an optical signal in a biased medium depends on both the light’s polarization and its direction of travel Lannebère et al. (2022); Rappoport et al. (2023); Shi et al. (2022); Morgado et al. (2024); Hakimi et al. (2024); de Sousa et al. (2024); Lannebère et al. (2025). This effect is tied to the Berry curvature of the system and has been shown to produce non-Hermitian (NH) electro-optic responses, leading to current-induced optical gain in both two- and three-dimensional materials Rappoport et al. (2023); Morgado et al. (2024). This phenomenon can also be explored in the context of lasing, demonstrating that NH effects can lead to polarization-dependent gain in biased systems Hakimi et al. (2024); Morgado et al. (2024); Lannebère et al. (2025). These works highlight the role of Berry curvature dipoles in shaping electro-optic effects, particularly in materials with low symmetry.
@@
To perform the symmetry analysis, we investigate the allowed components of the Berry curvature dipole tensor with the tools implemented in the Bilbao Crystallographic Server (BCS) Aroyo et al. (2006a, b). We explore various symmetries and examine different electric bias configurations. Notably, we demonstrate that the non-Hermitian electro-optic effect can be engineered to produce optical gain for all polarizations of the optical field (in the plane perpendicular to the bias), which can be switched to optical dissipation simply by reversing the orientation of the static electric bias. Our analysis also reveals that gain can occur in the absence of any gyrotropic effect.
Finally, we illustrate the practical implications of this effect by examining the reflectance of materials in setups involving mirrors. This allows us to evaluate how the incoming wave polarization can trigger either dissipative or gain responses in systems with different symmetries, providing valuable insights into the potential applications of these materials in optical devices and electromagnetic systems.
@@ II Electro-optic permittivity
The linearized optical response of a generic low-symmetry three-dimensional metal under a static electric bias $\mathbf{E}_{0}$ can be found using the semiclassical Boltzmann transport theory Rappoport et al. (2023); Morgado et al. (2024). The linear electro-optic response is determined by the Berry curvature dipole tensor $\overline{{\bf D}}$,
$D_{ij}=-\frac{1}{(2\pi)^{3}}\int\frac{\partial f^{0}_{\mathbf{k}}}{\partial k_{i}}\Omega_{\mathbf{k},j}d^{3}{\mathbf{k}}=\frac{1}{(2\pi)^{3}}\int f^{0}_{\mathbf{k}}\frac{\partial{\Omega_{\mathbf{k},j}}}{\partial k_{i}}d^{3}{\mathbf{k}},$ (1)
@@ III Chiral-gain and Linear-dichroic gain
Table 1: Classification of the real and imaginary components of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime}$ and $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ based on their reciprocal and non-reciprocal, Hermitian and non-Hermitian properties.
@@
Here, it is implicit that the fields have a time-harmonic variation $\mathbf{E}(t)=\frac{1}{2}\left(\mathbf{E}_{\omega}e^{-i\omega t}+\mathbf{E}_{\omega}^{\ast}e^{i\omega t}\right)$. Thus, the quadratic form $\mathbf{E}_{\omega}^{\ast}\cdot\overline{{\bf\varepsilon}}^{\prime\prime}\cdot\mathbf{E}_{\omega}$ governs the direction of the irreversible energy flow. For a positive $\mathbf{E}_{\omega}^{\ast}\cdot\overline{{\bf\varepsilon}}^{\prime\prime}\cdot\mathbf{E}_{\omega}$ the energy of the wave is irreversibly dissipated in the material in the form of heat. Conversely, for a negative $\mathbf{E}_{\omega}^{\ast}\cdot\overline{{\bf\varepsilon}}^{\prime\prime}\cdot\mathbf{E}_{\omega}$ the medium supplies energy to the wave, corresponding to optical gain. The tensor $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ describes the non-Hermitian interactions that arise from the electro-optic effect. It combines additively with the standard linear (Drude) response contribution of the material Morgado et al. (2024); Rappoport et al. (2023).
@@
is by definition the chiral-gain vector. Thus, the light-matter interactions can be either dissipative ($p_{\rm d,EO}^{\rm I}>0$) or active ($p_{\rm d,EO}^{\rm I}<0$), depending on the relative alignment of the spin angular momentum of the wave and the chiral-gain vector. The optimal polarization to unlock the gain is a circular polarization in the plane perpendicular to the chiral-gain vector, whereas the optimal polarization to unlock the dissipative properties of the medium is a circular polarization in the same plane but with opposite handedness Lannebère et al. (2025). Accordingly, the two eigenvectors of $i\overline{{\bf\varepsilon}}_{\text{EO,I}}^{\prime\prime}$ are two circular polarizations in the plane perpendicular to ${\bf{\Omega}}_{\omega}$. As shown in table 1, $i\overline{{\bf\varepsilon}}_{\text{EO,I}}^{\prime\prime}$ also contributes to a nonreciprocal electromagnetic response. Note also that as the gyrotropic and chiral-gain terms have the same mathematical structure, they typically coexist within the same material.
@@ IV Symmetry Analysis of the Non-Hermitian Response
Figure 1: Classification diagram based on the work of Halasyamani and Poeppelmeier Halasyamani and Poeppelmeier (1998), dividing non-centrosymmetric materials into seven categories according to their symmetry-dependent properties. Only the five categories that can have a non-zero Berry curvature dipole are shown in the diagram. The diagram classifies materials by separating them into polar and nonpolar crystal sets and also organizing them based on properties such as optical activity, circular dichroism, and piezoelectricity. The diagram also serves to categorize the Berry curvature dipole tensor of various materials.
@@
In this article, we will focus on specific symmetry categories shown in the diagram in Figure 1 that are distinct from those examined in our previous studies Lannebère et al. (2022); Rappoport et al. (2023); Morgado et al. (2024). Furthermore, we will focus on point groups that possess a principal axis, with the static electric bias aligned parallel to this axis. For materials where the principal axis is aligned with the $z$-direction, the BD tensor assumes the generic form:
@@
Consistent with the analysis in Sect. III, the tensor $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ has a trivial eigenvalue ($\lambda_{3}=0$). Consequently, the non-Hermitian electro-optic response is determined by two orthogonal eigenpolarizations, ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$, which lie in the $xoy$ plane, perpendicular to the bias field ${\bf{E}}_{0}$. Notably, the chiral-gain component of the response appears only for point groups where $D_{xx}+D_{yy}$ can be nontrivial. More generally, the trace of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$, which equals the sum of the eigenvalues, is given by $D_{yx}-D_{xy}$. Therefore, when $D_{yx}=D_{xy}$, the response becomes indefinite, and $\lambda_{2}=-\lambda_{1}$.
@@ IV.1 Pure linear dichroic gain: Categories B, C and E
One of the most interesting scenarios originates from the BD of the group $mm2$, which belongs to category B. For this group, considering a polar axis directed along $z$, the Berry curvature dipole has the form in Eq. (14) with all diagonal elements equal to zero.
@@
Figure 2: Eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of the non-Hermitian part of the electro-optic permittivity tensor $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ for an electric bias along the principal axis and different point group symmetries belonging to the different categories. Negative eigenvalues indicate gain whereas positive eigenvalues correspond to loss. The insets illustrate the eigenpolarizations of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$. The numerical parameters are $D_{xy}=1$, $D_{0}=1.5$, $\tau=1\,\rm{ps}$ and $\omega_{0z}/{(2\pi)}\approx 66.47\,\rm{THz}$ ($E_{0}=10^{4}$ V/m).
Point groups belonging to category C represent a special case of the BD tensor for the group $mm2$, where the BD is purely antisymmetric $D_{yx}=-D_{xy}$. These materials are polar but do not present any optical activity. The nontrivial eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ are doubly degenerate and given by
@@
This response underscores the critical importance of the symmetry properties of the material and their impact on the non-Hermitian electro-optic response of the system. When $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ is real and symmetric, the system has a reciprocal (non-gyrotropic) response with optical gain. This feature is particularly relevant for applications requiring polarization-independent optical gain, while still allowing the gain to be controlled via the applied bias.
It is emphasized that when an electric bias is applied along the polar axis, the discussed groups do not exhibit any electro-optic-induced nonreciprocal gyrotropic response. This is because, as previously noted, a gyrotropic response always coexists with a chiral-gain response, which is absent in the considered point groups.
@@
Hence, these materials may behave as indefinite gain media with the gain and dissipation activated by orthogonal linear polarizations. The $\overline{4}2m$ point group represents a particular case of the $\overline{4}$ group where $D_{xy}=0$. In this case, the two eigenvalues display a frequency dependence that is symmetric about the frequency axis, as illustrated by the green curves in Fig. 2.
@@ IV.4 Biases along other crystallographic directions
For a horizontal bias (relative to the symmetry axis), it can be shown that only the elements $\varepsilon_{xz}^{\prime\prime}$, $\varepsilon_{yz}^{\prime\prime}$, $\varepsilon_{zx}^{\prime\prime}$, and $\varepsilon_{zy}^{\prime\prime}$ are nonzero. In particular, the trace of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ is zero, implying that the sum of all eigenvalues must also be zero. Moreover, the determinant of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ is identically zero for this type of structure, meaning that one of its eigenvalues necessarily vanishes, while the remaining two are equal in magnitude but opposite in sign. In general, the $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ has both chiral-gain and linear-dichroic gain components, resulting in elliptical eigenpolarizations that depend on frequency.

### v2

@@ Abstract
Here, we investigate how crystal symmetry tailors the non-Hermitian electro-optic effect arising from the Berry curvature dipole. Specifically, we demonstrate the critical influence of the material’s point group symmetry and external electric biases in shaping this effect, leading to current-induced optical gain and non-reciprocal optical responses. Through a symmetry-based analysis of the crystallographic point groups, we identify how different symmetries affect the electro-optic response, enabling the engineering of polarization-dependent optical gain without the need for gyrotropic effects. In particular, we demonstrate that the non-Hermitian electro-optic response in a broad class of crystals is characterized by linear dichroic gain. In this type of response, the eigenpolarizations that activate the gain or dissipation are linearly polarized. Depending on the point group symmetry, it is possible to achieve gain (or dissipation) for all eigenpolarizations or to observe polarization-dependent gain and dissipation. Weyl semimetals emerge as promising candidates for realizing significant non-Hermitian electro-optic effects and linear dichroic gain. We further examine practical applications by studying the reflectance of biased materials in setups involving mirrors, demonstrating how optical gain and attenuation can be controlled via symmetry and bias configurations.
@@ I Introduction
Several studies have further linked the BD to optical effects, including SHG, the kinetic Faraday effect, where the rotatory power is proportional to the electric current and reverses sign with the applied electric field Vorob’ev et al. (1979); Tsirkin et al. (2018); König et al. (2019), and the circular photogalvanic effect, where the photocurrent reverses sign with the helicity of light Ivchenko and Pikus (1978); Asnin et al. (1978); Deyo et al. (2009); Xu et al. (2018). These effects are particularly prominent in materials with small band gaps, where the Berry curvature gets resonantly enhanced and significantly influences interband transitions. Crucially, many second-order optical responses have been shown to directly correlate with the Berry curvature dipole, establishing them as effective probes for investigating a material’s electronic structure and symmetry Xu et al. (2018); Bhalla et al. (2022).
Recent proposals have demonstrated that electro-optic effects can lead to non-reciprocal optical gain, where the amplification of an optical signal in a biased medium depends on both the light’s polarization and its direction of travel Lannebère et al. (2022); Rappoport et al. (2023); Shi et al. (2022); Morgado et al. (2024); Hakimi et al. (2024); de Sousa et al. (2024); Lannebère et al. (2025); Ma et al. (2025). This effect is tied to the Berry curvature of the system and has been shown to produce non-Hermitian (NH) electro-optic responses, leading to current-induced optical gain in both two- and three-dimensional materials Rappoport et al. (2023); Morgado et al. (2024). This phenomenon can also be explored in the context of lasing, demonstrating that NH effects can lead to polarization-dependent gain in biased systems Hakimi et al. (2024); Morgado et al. (2024); Lannebère et al. (2025). These works highlight the role of Berry curvature dipoles in shaping electro-optic effects, particularly in materials with low symmetry.
@@
To perform the symmetry analysis, we investigate the allowed components of the Berry curvature dipole tensor with the tools implemented in the Bilbao Crystallographic Server (BCS) Aroyo et al. (2006a, b). We explore various symmetries and examine different electric bias configurations. Notably, we demonstrate that the non-Hermitian electro-optic effect can be engineered to produce optical gain for all polarizations of the optical field (in the plane perpendicular to the bias), which can be switched to optical dissipation simply by reversing the orientation of the static electric bias. Our analysis also reveals that gain can occur in the absence of any gyrotropic effect (optical rotation).
Finally, we illustrate the practical implications of the NHEO effect by examining the reflectance of materials in setups involving mirrors. This allows us to evaluate how the incoming wave polarization can trigger either dissipative or gain responses in systems with different symmetries, providing valuable insights into the potential applications of these materials in optical devices and electromagnetic systems.
@@ II Electro-optic permittivity
The linearized optical response of a generic low-symmetry three-dimensional conductor under a static electric bias $\mathbf{E}_{0}$ can be found using the semiclassical Boltzmann transport theory Rappoport et al. (2023); Morgado et al. (2024). The linear electro-optic response is determined by the Berry curvature dipole tensor $\overline{{\bf D}}$,
$D_{ij}=-\frac{1}{(2\pi)^{3}}\int\frac{\partial f^{0}_{\mathbf{k}}}{\partial k_{i}}\Omega_{\mathbf{k},j}d^{3}{k}=\frac{1}{(2\pi)^{3}}\int f^{0}_{\mathbf{k}}\frac{\partial{\Omega_{\mathbf{k},j}}}{\partial k_{i}}d^{3}{k},$ (1)
@@ III Chiral-gain and Linear-dichroic gain
Table 1: Classification of the real and imaginary components of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime}$ and $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ based on their reciprocal and non-reciprocal, Hermitian and non-Hermitian properties. The term chiral gain refers to circular-dichroic gain.
@@
Here, it is implicit that the fields have a time-harmonic variation $\mathbf{E}(t)=\frac{1}{2}\left(\mathbf{E}_{\omega}e^{-i\omega t}+\mathbf{E}_{\omega}^{\ast}e^{i\omega t}\right)$. Thus, the quadratic form $\mathbf{E}_{\omega}^{\ast}\cdot\overline{{\bf\varepsilon}}^{\prime\prime}\cdot\mathbf{E}_{\omega}$ governs the direction of the irreversible energy flow. For a positive $\mathbf{E}_{\omega}^{\ast}\cdot\overline{{\bf\varepsilon}}^{\prime\prime}\cdot\mathbf{E}_{\omega}$ the energy of the wave is irreversibly dissipated in the material in the form of heat. Conversely, for a negative $\mathbf{E}_{\omega}^{\ast}\cdot\overline{{\bf\varepsilon}}^{\prime\prime}\cdot\mathbf{E}_{\omega}$ the medium supplies energy to the wave, corresponding to optical gain. The tensor $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ describes the non-Hermitian interactions that arise from the electro-optic effect. It combines additively with the standard linear (Drude) response contribution of the material Morgado et al. (2024); Rappoport et al. (2023), which is strictly dissipative.
@@
is by definition the chiral-gain vector. Thus, the light-matter interactions can be either dissipative ($p_{\rm d,EO}^{\rm I}>0$) or active ($p_{\rm d,EO}^{\rm I}<0$), depending on the relative alignment between the spin angular momentum of the wave and the chiral-gain vector. The optimal polarization to unlock the gain is a circular polarization in the plane perpendicular to the chiral-gain vector, whereas the optimal polarization to unlock the dissipative properties of the medium is a circular polarization in the same plane but with opposite handedness Lannebère et al. (2025). Accordingly, the two eigenvectors of $i\overline{{\bf\varepsilon}}_{\text{EO,I}}^{\prime\prime}$ are two circular polarizations in the plane perpendicular to ${\bf{\Omega}}_{\omega}$. As shown in table 1, $i\overline{{\bf\varepsilon}}_{\text{EO,I}}^{\prime\prime}$ also contributes to a nonreciprocal electromagnetic response. Note also that as the gyrotropic and chiral-gain terms have the same mathematical structure, they typically coexist within the same material.
@@ IV Symmetry Analysis of the Non-Hermitian Response
Figure 1: Classification diagram based on the work of Halasyamani and Poeppelmeier Halasyamani and Poeppelmeier (1998), dividing non-centrosymmetric materials into seven categories according to their symmetry-dependent properties. Only the five categories that can have a non-zero Berry curvature dipole are shown in the diagram. The diagram classifies materials by separating them into polar and nonpolar crystal sets, and also organizing them based on properties such as optical activity, circular dichroism, and piezoelectricity. The diagram also serves to categorize the Berry curvature dipole tensor of various materials.
@@
In this article, we will focus on specific symmetry categories shown in the diagram of Figure 1 that are distinct from those examined in our previous studies Lannebère et al. (2022); Rappoport et al. (2023); Morgado et al. (2024). Furthermore, we will focus on point groups that possess a principal axis, with the static electric bias aligned parallel to this axis. Assuming that the principal axis is oriented along the $z$-direction, the BD tensor takes the generic form:
@@
Consistent with the analysis in Sec. III, the tensor $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ has a trivial eigenvalue ($\lambda_{3}=0$). Consequently, the non-Hermitian electro-optic response is determined by two orthogonal eigenpolarizations, ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$, which lie in the $xoy$ plane, perpendicular to the bias field ${\bf{E}}_{0}$. Notably, the chiral-gain component of the response appears only for point groups where $D_{xx}+D_{yy}$ can be nontrivial. Furthermore, the trace of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$, which equals the sum of the eigenvalues, is given by $D_{yx}-D_{xy}$. Therefore, when $D_{yx}=D_{xy}$, the response becomes indefinite, and $\lambda_{2}=-\lambda_{1}$.
@@ IV.1 Pure linear dichroic gain: categories B, C and E
One of the most interesting scenarios originates from the BD of the point group $mm2$, which belongs to category B. For this group, considering a polar axis directed along $z$, the Berry curvature dipole has the form in Eq. (14) with all diagonal elements equal to zero.
@@
Figure 2: Eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of the non-Hermitian part of the electro-optic permittivity tensor $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ for an electric bias along the principal axis and different point group symmetries belonging to the different categories (the category is indicated by a letter in parentheses). Negative eigenvalues indicate gain whereas positive eigenvalues correspond to loss. The insets illustrate the eigenpolarizations of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$. The numerical parameters are $D_{xy}=1$, $D_{0}=1.5$, $\tau=1\,\rm{ps}$ and $\omega_{0z}/{(2\pi)}\approx 66.47\,\rm{THz}$ ($E_{0}=10^{4}$ V/m).
Point groups belonging to category C represent a special case of the BD tensor for the group $mm2$, where the BD is purely antisymmetric: $D_{yx}=-D_{xy}$. These materials are polar but do not present any optical activity. The nontrivial eigenvalues $\lambda_{1}$ and $\lambda_{2}$ of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ are doubly degenerate and given by
@@
This response underscores the critical importance of the symmetry properties of the material and their impact on the non-Hermitian electro-optic response. When $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ is real and symmetric, the system has a reciprocal (non-gyrotropic) response with optical gain. This feature is particularly relevant for applications requiring polarization-independent optical gain, while still allowing the gain to be controlled via the applied bias.
We emphasize that when an electric bias is applied along the polar axis, the discussed groups do not exhibit any electro-optic-induced nonreciprocal gyrotropic response. This is because, as previously noted, a gyrotropic response always coexists with a chiral-gain response, which is absent in the considered point groups.
@@
Hence, these materials may behave as indefinite gain media with the gain and dissipation activated by orthogonal linear polarizations. The response in the $\overline{4}2m$ point group represents a particular case of that in the $\overline{4}$ group where $D_{xy}=0$. In this case, the two eigenvalues display a frequency dependence that is symmetric about the frequency axis, as illustrated by the green curves in Fig. 2.
@@ IV.4 Biases along other crystallographic directions
For a horizontal bias (relative to the symmetry axis), it can be shown that only the elements $\varepsilon_{xz}^{\prime\prime}$, $\varepsilon_{yz}^{\prime\prime}$, $\varepsilon_{zx}^{\prime\prime}$, and $\varepsilon_{zy}^{\prime\prime}$ are nonzero. In particular, the trace of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ is zero, implying that the sum of all eigenvalues must also be zero. Moreover, the determinant of $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ is identically zero for this type of matrix structure, meaning that one of its eigenvalues necessarily vanishes, while the remaining two are equal in magnitude but opposite in sign. In general, the $\overline{{\bf\varepsilon}}_{\text{EO}}^{\prime\prime}$ has both chiral-gain and linear-dichroic gain components, resulting in elliptical eigenpolarizations that depend on frequency.